OpenStudy (anonymous):

ln(x)*ln(x-1) Suppose you wanted to prep this fcn for L'Hopital's Rule. The expression is the limit as x approaches 1 from the right side of the above fcn. How do we do so? ln(x)*ln(x-1) Suppose you wanted to prep this fcn for L'Hopital's Rule. The expression is the limit as x approaches 1 from the right side of the above fcn. How do we do so? @Mathematics

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0+}\ln(x)\ln(x-1)\]

OpenStudy (anonymous):

forget it

OpenStudy (anonymous):

you cannot take the limit as x approaches 0 of ln(x-1)

OpenStudy (anonymous):

domain \[\ln(x-1)\] is \[x>1\]

OpenStudy (anonymous):

and if it is \[\lim_{x\rightarrow 1^+}\ln(x)\ln(x-1)\] l\hopital would not apply because this is of the form \[1\times -\infty\]

OpenStudy (phi):

I think he made a typo. He wants the expression is the limit as x approaches 1 from the right side of the above fcn.

OpenStudy (anonymous):

ok then it is not "undetermined" or whatever they say

OpenStudy (anonymous):

limx->0 ln(x).ln(x-1) = limx->0 ln(x)./1/ln(x-1) <- use L'Hopital's rule now limx->0 1/x/1/1/(x-1) = limx-> (x-1)/x <-use L'Hopital's rule again limx->0 (1)/1) = 1

OpenStudy (anonymous):

How do we make it an indeterminate form?

OpenStudy (anonymous):

you would make a mistake to write it as a fraction say, and use l'hoptital because that only works for undetermined limits. like for example \[\frac{0}{0}\] or \[\frac{\infty}{\infty}\] or \[1^{\infty}\] etc

OpenStudy (anonymous):

ln(x) / 1/ln(x-1) yields 0/infinity, and this isn't in L'H form...

OpenStudy (anonymous):

hold the phone you do not " make it an indeterminate form"

OpenStudy (anonymous):

it either is in one or it is not to begin with

OpenStudy (anonymous):

So what the heckk do I do for this case?

OpenStudy (anonymous):

lets go slow.

OpenStudy (zarkon):

what is the actual problem?

OpenStudy (anonymous):

QRAwarrior, check my answer again :)

OpenStudy (anonymous):

if i want the limit as x goes to 0 of 1/x, i do not "change the form" that limit does not exist, so i cannot make it exist by changing the question

OpenStudy (anonymous):

The problem is this: Limit as x approaches 0 from the right side of (lnx)(ln(x-1))

OpenStudy (zarkon):

are you sure

OpenStudy (zarkon):

is this a course in complex analysis?

OpenStudy (anonymous):

send me a medal later :) brb

OpenStudy (anonymous):

what you are thinking of is how to change something like \[x\ln(x)\] into \[\frac{\ln(x)}{\frac{1}{x}}\]so that i can turn an indermined limit like \[\lim_{x\rightarrow 0}x\ln(x)\] which is of the form \[0\times -\infty\] into \[\frac{\ln(x)}{\frac{1}{x}}\] which is of the form \[\lim_{x\rightarrow 0^+}\frac{-\infty}{\infty}\] that is just algebra

OpenStudy (anonymous):

but if your limit is not undetermined to begin with, you cannot change it into something that is

OpenStudy (anonymous):

maybe i am not sure of the actual question, so let me be quiet

OpenStudy (phi):

@Q, you keep saying x->0 but the original post says as x->1 (which makes much more sense) for lim x->1 we know ln(x) approaches 0, so we might re-write the expression as ln(x-1)/ (1/ln(x))

OpenStudy (anonymous):

ooooooooooh ok that you can do. same gimmick as before as phi wrote